### v1

# RSTeller: Scaling Up Visual Language Modeling in Remote Sensing with Rich Linguistic Semantics from Openly Available Data and Large Language Models

Abstract
Abundant, well-annotated multimodal data in remote sensing are pivotal for aligning complex visual remote sensing (RS) scenes with human language, enabling the development of specialized vision language models across diverse RS interpretation tasks. However, annotating RS images with rich linguistic semantics at scale demands expertise in RS and substantial human labor, making it costly and often impractical. In this study, we propose a workflow that leverages large language models (LLMs) to generate multimodal datasets with semantically rich captions at scale from plain OpenStreetMap (OSM) data for images sourced from the Google Earth Engine (GEE) platform. This approach facilitates the generation of paired remote sensing data and can be readily scaled up using openly available data. Within this framework, we present RSTeller, a multimodal dataset comprising over 1 million RS images, each accompanied by multiple descriptive captions. Extensive experiments demonstrate that RSTeller enhances the performance of multiple existing vision language models for RS scene understanding through continual pre-training. Our methodology significantly reduces the manual effort and expertise needed for annotating remote sensing imagery while democratizing access to high-quality annotated data. This advancement fosters progress in visual language modeling and encourages broader participation in remote sensing research and applications. The RSTeller dataset is available at https://github.com/SlytherinGe/RSTeller.
keywords:
Vision language model , Multimodal dataset , OpenStreetMap , Google Earth Engine , Large language models
journal: ISPRS Journal of Photogrammetry and Remote Sensing
1 Introduction
Vision language models (VLMs) hold great promise in remote sensing (RS) as they offer a means to interpret complex visual scenes for non-expert users in a human-like manner and support the development of diverse RS applications. However, the effective training of VLMs hinges on the availability of large quantities of paired data to ensure the alignment of visual and linguistic semantics for optimal performance. The substantial data requirements present a significant challenge in scaling up VLMs for RS, impeding their broader applicability and performance enhancement. Consequently, there is a pressing need for automated, scalable, and efficient methods to generate datasets with paired RS imagery and semantic-rich captions. Such methods are essential for overcoming the bottleneck associated with data scarcity and facilitating the development of robust VLMs tailored for RS tasks. Moreover, the availability of an out-of-the-box, well-annotated dataset comprising RS images with rich linguistic semantics holds immense value. It not only streamlines the initial training of VLMs but also contributes to reducing the environmental impact associated with the individual generation of such data, thereby addressing sustainability concerns.
1.1 Related works
The intersection of computer vision and natural language processing (NLP) has witnessed remarkable advancements in recent years, propelled by the emergence of large-scale pre-trained models. When CLIP (Radford et al., 2021) was initially introduced, it swiftly captivated the research community with its remarkable performance in zero-shot classification and cross-modal retrieval tasks, rivaling some state-of-the-art models trained with full supervision. CLIP achieves this feat by aligning image and text representations through a contrastive learning strategy, endowing it with the capability for zero-shot inference using the learned representations. Inspired by this paradigm, the RS community embarked on leveraging similar strategies to construct VLMs tailored for RS tasks. Notable efforts such as those by Liu et al. (2024a), Zhang et al. (2023), and Wang et al. (2024) have undertaken continual pre-training of CLIP on paired RS data, thereby extending its capabilities to encompass RS scene understanding.
Concurrently, the advent of large language models (LLMs), exemplified by ChatGPT (Brown et al., 2020; Ouyang et al., 2022), alongside other models (Brown et al., 2020; Ouyang et al., 2022; Du et al., 2021; Jiang et al., 2024), has significantly enriched the landscape of NLP. These LLMs possess the remarkable ability to interact with humans through natural language, facilitate in-context learning, instruction following (Ouyang et al., 2022), and reasoning (Wei et al., 2022), showcasing their versatility across diverse linguistic tasks. Motivated by the power of LLMs, researchers have proposed to align visual models with large language models to achieve the SOTA performance on image captioning (Li et al., 2022), visual question answering (VQA) (Li et al., 2023a; Zhu et al., 2023; Dai et al., 2024; Liu et al., 2024b), visual grounding (Wang et al., 2023b), and etc. In the RS domain, efforts have mirrored these trends. For instance, RSGPT (Hu et al., 2023) fine-tuned InstructBLIP (Dai et al., 2024) on a self-built dataset to achieve state-of-the-art performance on RS image captioning and VQA tasks. Remote Sensing ChatGPT (Guo et al., 2024) utilizes ChatGPT to interpret user queries, plan remote sensing interpretation tasks, and generate final responses for users. Additionally, GeoChat (Kuckreja et al., 2024) fine-tuned LLaVA-1.5 (Liu et al., 2024b) using LoRA (Hu et al., 2022), enabling its support for multi-task and multi-round conversion with users to interact with RS imagery.
Despite the progress in developing vision language models tailored for RS tasks, a significant gap remains between these models and those designed for general purposes. One of the main differences between the two paradigms lies in the scale of the models and their training data. Several studies have empirically observed that model performance is positively correlated with the model size and the data scale (Kaplan et al., 2020; Riquelme et al., 2021; Tay et al., 2021; Zhai et al., 2022; Cherti et al., 2023). Notably, the CLIP study trained its largest model, ViT-L/14, on a private WebImageText (WIT) dataset containing 400 million image-text pairs, resulting in a substantial leap in performance compared to models trained on smaller datasets. The success of CLIP has motivated the development of subsequent VLMs with increased model size and data scale for better performance. For instance, ALIGN (Jia et al., 2021) trained a large EfficientNet-L2 image encoder on a private dataset of 1.8 billion image-text pairs. CoCa (Yu et al., 2022) used the ALIGN dataset as well as a private dataset called JFT-3B, containing 3 billion image-text pairs, to train a ViT-g/14 visual encoder and a language model. PaLI (Chen et al., 2022) further expanded this approach by training a multi-language, multi-task VLM using ViT-e (4B parameters) and mT5-XXL (13B parameters) on a private dataset of 29 billion image-text pairs.
The ever-growing model sizes and their data demands have quickly surpassed the capacity of earlier open image-text datasets like Microsoft COCO (Lin et al., 2014), Visual Genome (Krishna et al., 2017), YFCC100M (Thomee et al., 2016), and CC3M (Sharma et al., 2018), which are insufficient for researchers without access to large-scale private datasets. Recently, the availability of large-scale multimodal data via the Internet has led to the creation of even larger image-text datasets, such as LAION-400M (Schuhmann et al., 2021), COYO-700M (Byeon et al., 2022), and LAION-5B (Schuhmann et al., 2022). These datasets have significantly enriched the VLM research community, enabling the development of many advanced VLMs (Li et al., 2023a; Wang et al., 2023b; Zhu et al., 2023).
Fig. 1: Comparison of existing image-text multimodal RS datasets. The Measure of Textual Lexical Diversity (MTLD) scores quantify the semantic richness of the captions. All captions were tokenized using the Natural Language Toolkit (NLTK) word tokenizer, and the color of each point indicates the scale of the tokens. The proposed RSTeller dataset is marked with a star. Its MTLD score is more than twice that of the next most diverse dataset, while also being substantial in data size.
The RS community has also made significant efforts to address the data scarcity problem for training VLMs for RS tasks. For readers’ easy comprehension of the scales of the related works, we compared them as in Fig. 1. Before the rise of large-scale pre-trained multimodal models like CLIP, several image-text datasets were developed, including UCM-Captions (Qu et al., 2016), Sydney-Captions (Qu et al., 2016), RSICD (Lu et al., 2017), RSITMD (Yuan et al., 2022), NWPU-Captions (Cheng et al., 2022), and RSVGD (Zhan et al., 2023). Among these, the largest dataset, NWPU-Captions, contains 31,500 images and 157,500 image-text pairs. The limited size of these datasets poses challenges for training VLMs like CLIP to perform well on RS tasks. To address this, (Liu et al., 2024a) compiled a RemoteCLIP dataset from three different types of RS tasks (segmentation, detection, and retrieval) across 17 individual datasets. They used hand-crafted rules to convert segmentation and detection annotations into multiple linguistic captions, resulting in a dataset with 165,745 images and 828,725 image-text pairs. Additionally, Zhang et al. (2023) proposed leveraging existing large-scale image-text datasets by filtering data related to RS and captioning several RS classification datasets with a pre-trained image caption model, creating an RS-related multimodal dataset called RS5M with 5 million image-text pairs. Training VLMs on these larger RS datasets has demonstrated improvements in their performance. However, there are limitations in building datasets from existing ones, as the upper bound is restricted by the quantity of available well-crafted datasets, which is limited compared to the vast Internet resources. In contrast, datasets like LAION (Schuhmann et al., 2021, 2022) were collected at the Internet scale, offering significantly larger amounts of data.
Despite the shortage of direct RS-related image-text datasets, there is an abundance of publicly available RS data. The Google Earth Engine (GEE) platform [^1] https://developers.google.com/earth-engine (Gorelick et al., 2017) provides a rich source of RS imagery, with comprehensive coverage of the Earth’s surface and access to a diverse range of aerial imagery from multiple sources. These sources vary in ground sampling distances, spectral bands, and sampling times, making them ideal for large-scale imagery data collection. On the other hand, OpenStreetMap (OSM) [^2] https://wiki.openstreetmap.org (Haklay and Weber, 2008) is a collaborative project that collects geo-elements worldwide, providing a rich source of geo-referenced data. OSM serves as a semantic source where geo-elements are annotated with descriptive tags, each composed of a key-value pair. The key describes a topic, category, or feature type (e.g., highway or name), while the value provides details for the specified feature (e.g., name: Jeff Memorial Highway). OSM currently contains over 96,000 distinct keys and over 155 million distinct tags, making it a valuable resource for harvesting geo-related linguistic data. Recently, Wang et al. (2024) proposed using GEE and OSM as data sources to build a large-scale image-text dataset for RS called SkyScript, comprising 2.6 million image-text pairs. They selected valid OSM elements for given RS images and designed a heuristic caption-assembling method to convert tags into captions. The dataset exhibits great generality for downstream tasks when training a CLIP model and shows scalable potential since it uses openly available data. However, the captions in the dataset are all simple phrases rather than semantically rich sentences, as they are merely combinations of the tag words. This limitation affects the semantic richness of the dataset and its further applicability.
Using LLMs to annotate data has become increasingly common as the performance of these methods improves. For instance, Li et al. (2023b) proposed using multiple VLMs and ChatGPT to refine the captions in image-text datasets, generating detailed image captions. To address the scarcity of well-annotated multimodal data, Wang et al. (2023a) introduced Caption Anything (CAT), which utilizes the Segment Anything Model (SAM) (Kirillov et al., 2023) and ChatGPT. CAT supports diverse visual and language controls to produce controllable captions. In another approach, Fan et al. (2024) used LLaMa-7B (Touvron et al., 2023) to rewrite the captions in the training set of a CLIP model as a method of data augmentation, and validated the effectiveness of these rewrites in the downstream tasks. In the field of RS, GeoChat (Kuckreja et al., 2024) employed Vicuna-1.5-7B (Chiang et al., 2023) to generate multi-round question and answer pairs for existing labeled RS datasets focused on object detection, visual question answering, or scene classification tasks. This resulted in a dataset with 306,000 image-instruction pairs, enabling the VLM to follow human instructions and handle multiple RS-related tasks. These methods demonstrate the feasibility and value of generating multimodal data with LLMs, inspiring us to design a large-scale multimodal dataset for RS using LLMs.
1.2 Contributions
In this work, we leverage openly available remote sensing data as an abundant resource for multimodal data generation. Specifically, we utilize the GEE platform to collect unlabeled aerial images, and the OSM platform as our semantic source. We employ Mixtral-7B (Jiang et al., 2024), an open-source LLM, to interpret plain OSM data. The OSM tag contents and their combinations are used to generate fluent, semantically rich descriptions for the aerial images. To ensure reproducibility and facilitate further scaling up of a dataset suitable for VLM training, we develop a comprehensive dataset generation workflow that encompasses raw data fetching, LLM-based captioning, and final dataset compilation.
Through this workflow, we introduce RSTeller, a multimodal dataset consisting of over one million remote-sensing images, each accompanied by multiple descriptive captions. We use the Measure of Textual Lexical Diversity (MTLD) score (McCarthy and Jarvis, 2010) as an indicator of the semantic richness of the captions. Our dataset is over two times more semantically rich than existing datasets (Fig. 1). Given the computational demands of LLM inference, our dataset provides a ready-to-use resource for researchers, eliminating the need for individual data generation.
Extensive experiments were conducted to validate the proposed methods and provide valuable insights for researchers training VLMs on RS data. First, we studied the relationship between the performance of the VLMs and the scale of RS domain data, highlighting the potential of further scaling up the training data. Then, we evaluated the effectiveness of RSTeller by continuing the pre-training of multiple VLMs and testing them on downstream tasks. The proposed data generation workflow and dataset are expected to be instrumental in advancing more sophisticated VLMs within the remote sensing field.
1.3 Paper structure
The structure of this paper is as follows: Section 2 details the proposed automatic data generation workflow. Section 3 describes and discusses the generated dataset, RSTeller, exploring its attributes in detail. Section 4 presents extensive experiments that explore the model performance concerning the data scale and the effectiveness of the proposed dataset. Section 5 provides some findings and suggestions according to our experiments. Section 6 outlines the limitations of the proposed work. Finally, Section 7 offers a summary of the research.
2 Automated Workflow for Dataset Generation
Fig. 2: Workflow for dataset generation. The left section outlines the main processes: data fetching, LLM-based captioning, and final dataset compilation. The right section details the specific captioning tasks that enhance the quality and richness of the generated captions.
In this section, we detail the automated workflow developed to generate a large-scale, RS-related multimodal dataset from openly available data. The overall workflow can be divided into two parts and is shown in Fig. 2. The left panel of the figure illustrates the main processes, including data fetching, LLM-based captioning, and final dataset compilation. All the processes can be run in parallel for the maximum efficiency. The right panel provides details of the captioning tasks that facilitate the generation of semantically rich captions.
2.1 Main Processes
2.1.1 Raw Data Acquisition
The first process in the workflow is to fetch raw data from the GEE and OSM platforms into well-defined local databases. For ease of reference, we designate these databases as the ”raw data database,” as depicted in Process 1 of Fig. 2. This database is crucial for preserving the downloaded data and enabling asynchronous handling of different processes to accelerate the entire workflow. The GEE platform provides a vast collection of aerial images, each typically covering a wide range of geographical locations with tens or hundreds of thousands of pixels. The size of these images is too large for direct use in training a VLM, necessitating the cropping of images into smaller patches. GEE offers several flexible APIs for patch extraction and data download, from ee.Image.getThumbURL for JPEG or PNG formats to ee.Image.getDownloadURL for GeoTIFF or NumPy formats. For detailed API usage instructions, the official documentation [^3] https://developers.google.com/earth-engine/apidocs is recommended.
To enhance file system performance, a database such as MongoDB is preferred and used in this work for storing the downloaded image patches, as a plain file system may be inefficient for large data indexing and searching. In addition to the imagery data, metadata such as patch coordinates and image timestamps are also required and carefully preserved for OSM data downloading and further analysis. An SQL-based database like MySQL or PostgreSQL is recommended for storing these metadata.
The OSM data is another important source of semantic information for remote sensing. It can be queried via the Overpass API [^4] https://wiki.openstreetmap.org/wiki/Overpass_API using specific OverpassQL queries. OSM contains three types of geo-elements: nodes, ways, and relations. As nodes are points and too small to be seen in an RS image, we focus only on ways and relations. We implemented a two-step method to query OSM data to ensure the inclusion of all necessary geo-elements. The first step is to query all the ways and relations that intersect with or are within the image patches. This step includes most necessary geo-elements but may miss some large geometries that do not intersect but enclose the image patches, such as lakes or forests. The second step is to query all the ways and relations that enclose the center point of the patch, including those large geo-elements missed in the first step but are still very important to the image content.
The geographical contents of the image patches may evolve over time. Since the OSM data is not updated in real-time, we queried the OSM data for each image patch up to one month after it was captured by the sensor to best reflect the current image content. However, querying from the official endpoint [^5] https://overpass-api.de/api/interpreter requires following a strict usage policy, allowing fewer than 10,000 queries and 1 GB of data per day. Some third-party backends like Kumi Systems Overpass Instance [^6] https://overpass.kumi.systems/api/interpreter provide a free usage policy but may not guarantee the capacity and availability for big data processing due to server overload and network traffic. In our experience, using the Kumi endpoint can fetch OSM data for around 100,000 patches, each covering about 72,253.44 square meters in a day. Considering there could be patches not covered by OSM data and that captioning valid patches consumes much more time, it is inefficient to query the remote server each time a patch needs captioning. Therefore, it is better to cache the OSM data in a local database ahead of time or deploy the Overpass server locally. Using local OSM data during captioning significantly reduces the query time and improves the overall efficiency.
2.1.2 Raw Caption Generation
Once the data fetching process is underway and a sufficient amount of raw data has been acquired, the workflow proceeds to the generation of raw captions. This process involves using an LLM, specifically Mixtral-7B (Jiang et al., 2024) in this work, to interpret and describe the OSM data associated with each image patch according to predefined tasks. Instead of using an instruction fune-tuned LLM, like Mixtral-7B-instruct, the use of the base LLM is advantageous. It provides greater flexibility in generating captions because it is not constrained by specific instructional patterns. This allows it for better following the instructions and examples provided by the prompt, making the tone and style of the generated captions more consistent with the intended purpose.
In this process, a patch that has not yet been captioned is first sampled from the raw data database. The corresponding OSM data is then fetched from the cached OSM database. The OSM data undergoes processing based on our predefined tasks, which will be discussed in Subsection 2.2. This step checks whether the OSM data contains tags that can be used to generate captions for the patch. Patches that contain no relevant way or relation elements other than administrative borders, or where the OSM elements are too small to be visualized, are deemed unusable. These unusable patches are discarded and labeled as such in the raw data database. Only OSM data from valid patches is retained for further processing.
The predefined tasks further interpret the attributes of the OSM elements, forming task-oriented prompts for the LLM. The LLM takes these prompts as input and generates captions for the image patches. Finally, the generated captions are saved in an annotation database, and the status of the image patches is updated in the raw data database to reflect their completion.
2.1.3 Caption Augmentation
Data augmentation in the linguistic aspect has been shown to enhance the training of vision language models (VLMs) (Fan et al., 2024). Unlike data augmentation in the visual domain, such as color jittering, noising, and random cropping, which can be easily implemented online during model training, augmenting the language aspect with consistent linguistic semantics and varied tones is resource-intensive, requiring the use of an LLM for high-quality results. Therefore, we incorporate a caption augmentation process into our workflow to generate multiple revisions of each caption.
During the caption augmentation process, captions that have not yet been revised are sampled from the annotation database. A specific revision task is defined to generate a prompt for the LLM based on the original caption and its initial task. The details of these tasks will be discussed in Subsection 2.2. The same LLM, Mixtral-7B, is used for this process. It takes the prompt as input and generates a revised caption for the image patch. The revised caption is then saved in the annotation database, and the status of the image patch is updated in the raw data database to reflect its completion.
This caption augmentation process is repeated until the desired number of revised captions is generated for each image patch. In RSTeller, our proposed dataset under this workflow, each image patch has at least two captions and up to five captions in total.
2.1.4 Dataset Compilation
As the annotation database populates and scales to the desired size, the final process in the workflow involves refining the annotations and compiling a dataset suitable for VLM training at scale.
During the inference of the LLM, some inevitable errors in the generated captions can lower the overall quality of the dataset. These errors can be categorized into two types: those that can be fixed and those that must be deleted. Fixable errors typically follow common patterns, such as repeated sentences within the same caption or undesired outputs appended after the caption ends, following some pattern of the input prompt. These redundant sentences and patterns can be removed through a heuristic process, allowing the remaining content to be used effectively. Unfixable errors usually include blank responses and duplicate captions for the same image patch. These errors cannot be corrected and must be deleted. The presence of these errors primarily stems from the imperfections of the LLM, and their frequency may be reduced through the evolution of the LLM, improvement of the prompts, or usage of a larger LLM.
Once the captions are refined, they are used to update the annotation database for future use, making the data ready for final compilation. Training a typical deep learning model on a given dataset usually involves randomly shuffling the data and sequentially iterating through the shuffled dataset. However, storing image patches and their annotations individually on the local disk and reading them randomly during training is time-consuming, especially when the dataset comprises millions of images. Additionally, training a VLM typically requires large batch sizes with multiple GPUs, resulting in an I/O bottleneck.
To address this issue and facilitate the distribution of the dataset over the internet, we propose compiling the dataset into multiple shards of tar files following the WebDataset format [^7] https://github.com/webdataset/webdataset. Each tar file comprises a set of image patches and their annotations. The RSTeller dataset, proposed in this work, is provided in this format. During training, multiple shards of the dataset can be prefetched into memory, then randomly shuffled and iterated through. This approach effectively improves I/O performance and results in better training efficiency.
2.2 Captioning Tasks
2.2.1 Descriptive Tasks
The objective of the descriptive task is to generate detailed and contextually accurate descriptions of the OSM data associated with each image patch. These descriptions should capture the essential features and elements present within the patch, such as geographical landmarks, infrastructure, vegetation, and other notable features. These descriptions serve as the raw captions and are produced in Process 2. In a single image patch, there usually are a number of OSM elements, each with multiple attributes and tags. The information provided by all the OSM elements in an image patch can be overwhelming, making it challenging for an LLM to generate a concise and accurate description.
To streamline the task and enhance the efficiency of LLM annotation, we propose focusing on one distinctive OSM element per image patch for the description. In OSM, a two-dimensional object bounded by one or more linear ways and marked by appropriate tags is recognized as an ”area.” Following this approach, we categorize all OSM elements in an image patch into two groups: ”area” and ”non-area.” The ”area” category includes OSM elements like land use, natural features, and buildings, while the ”non-area” category includes elements not recognized as an ”area,” such as roads, coastlines, and other linear features. Given the significant differences in geometry and attributes between these two categories, we have designed two distinct tasks: Task 1 for the ”area” category and Task 2 for the ”non-area” category. Fig. 7 presents examples from the RSTeller dataset, showcasing image patches, selected OSM elements, and corresponding captions. The selected OSM elements are highlighted in red on the image patches. Image patches (2), (3), and (13) are captioned based on the ”non-area” category, while all other image patches are captioned based on the ”area” category.
Task 1 (Area description): As is shown in Fig. 2, this task follows four steps: distinctive area selection, attribute interpretation, tag interpretation, and prompt assembly.
Distinctive area selection: In this step, the OSM element with the largest area within the image patch is selected as the most distinctive element to be described. The selected element undergoes a validation check to ensure it is not too small to be visible in the image. In this work, a threshold of 0.1 is used to filter out elements with an area smaller than this threshold portion of the image patch.
Attribute interpretation: To enrich the semantic information of the selected element and facilitate the annotation process of the LLM, we extract several predefined attributes of the selected element and interpret them using a heuristic rule-based method. We have defined five attributes to describe the element, as shown in Table 1, including coarse location, shape, normalized size, simplified geometry, and is-cropped. The coarse location attribute identifies the center of the element within a $3\times 3$ grid, assigning a label to each grid cell (e.g., ”left-top”, ”center”, ”right-bottom”). The shape attribute describes the approximate shape of the element based on its geometry, defined by four categories: ”square”, ”rectangular”, ”circular”, and ”irregular”. These categories are determined by hard thresholding on the element’s area-to-perimeter ratio and shape-to-area ratio. The normalized size attribute indicates the relative size of the element compared to the image patch. The simplified geometry attribute represents the simplified and normalized geometry of the element, formatted as coordinate pairs enclosed in brackets (x, y). To avoid overwhelming the LLM annotators with extremely complex geometries, we use the Douglas-Peucker algorithm (Ramer, 1972) to simplify the element’s geometry. As some complex elements may be represented by multiple polygons, we encapsulate the coordinate pairs of each polygon within a list, and all polygons of an element are enclosed within braces. In this work, (0, 0) represents the bottom-left corner and (1, 1) the top-right corner, with all coordinates normalized to the range of [0, 1]. The is-cropped attribute is a boolean value that indicates whether parts of the element extend beyond the image patch, with values ”True” or ”False”. This attribute is determined by checking if the element visible in the image patch is cropped from a larger geometry.
Table 1: Attributes for distinctive area description in captioning Task 1
Attribute Example
Coarse Location ”left-top”, ”center”, etc.
Shape ”square”, ”rectangular”, ”circular”, ”irregular”
Normalized Size ”0.300”, ”0.453”, etc.
Simplified Geometry ”{[(0.000, 0.500), (0.120, 0.400), …]}”, etc.
Is-cropped True, False
Tag interpretation: OSM tags can sometimes be simple yet ambiguous. For example, the tag ”man_made: works” might be confusing to both the LLM and humans, as it can refer to various man-made structures. The OSM tag Wiki database provides explanations for such tags. According to the tag Wiki, ”man_made: works” refers to a factory or industrial production plant. This clarification removes ambiguity and makes the tag more meaningful for the LLM. Therefore, we query all OSM tag descriptions of the selected element from the tag Wiki. Using these descriptions, we interpret the tags into a uniform and linguistically consistent format, instead of merely using the key-value pairs. For tags with unbounded values, such as ”name: Jeff Memorial Highway,” we focus on interpreting the tag key instead. This results in a more comprehensive and accurate description of the selected element. An example of tag interpretation of the two kinds is shown in Table 2.
Table 2: Examples of tag interpretation
Tag Interpretation
man_made: works man_made: works. The tag belongs to the tag group ”NULL”. This tag means: ”A factory or industrial production plant”.
name: Jeff Memorial Highway Its key is ”name”, which means ”the primary name: in general, the most prominent signposted name or the most common name in the local language(s).”. The tag belongs to a tag group ”names”. The tag value is Jeff Memorial Highway.
Prompt assembly: Once the attributes and tags have been interpreted, the final step in Task 1 is to assemble all the interpretations into a coherent prompt for the LLM. This involves organizing the interpreted attributes and tag descriptions into a structured format that the LLM can effectively process. In this work, we employ a template-based approach for prompt assembly, where we define a task-specific template to structure the prompt and fill in the placeholders with the interpreted attributes and tags. A brief showcase of the prompt template for Task 1 is presented in Fig. 3. The overall structure of the prompt is divided into three parts. First, general instructions are provided to the LLM under the headings ###Instruction### and ###Captioning Objective###. The former offers general task instructions and some basic knowledge about the attributes used, while the latter highlights the attributes crucial for the task and details that require particular attention. Specifically, we instruct the LLM to not only describe the selected element but also infer its possible surroundings and context based on the provided information. The inferred contexts by the LLM should be highlighted by some words denoting uncertainty, such as ”possible” or ”may”. This encourages the LLM to generate richer and more informative captions, expanding the scope of the task beyond the selected element to the entire image patch. Next, a one-shot example is included to illustrate the format and style of the desired output, helping the LLM understand the task requirements more clearly. The example follows the same format as the real task, consisting of a ”Raw” and ”Caption” pair. The ”Raw” part includes attributes and tags interpreted from an example image patch, while the ”Caption” part is the gold-standard caption from a human annotator and revised by GPT-4. Lastly, task-specific inputs are provided. The ”Raw” part is a predefined context with multiple placeholders, where the interpreted attributes and tags are directly filled, except for the ”Partition” attribute. If the ”Partition” attribute is True, the placeholder is replaced with the sentence ”Some parts of the geometry extended out of this ROI,” otherwise, it is left blank. The ”Caption” part remains blank, awaiting the LLM’s continual writing. This assembly step ensures that the LLM’s output is both precise and aligned with the intended use of the remote sensing data, facilitating effective and insightful captioning.
Fig. 3: A brief showcase of the prompt template for Task 1. Details are replaced by ellipses for brevity. Placeholders are wrapped in braces and are to be filled dynamically by the task.
Task 2 (Non-area description): Similar to Task 1, this task involves four steps. Based on our inspection of the collected OSM data, the ”non-area” with the longest length is sometimes a simple line from a large geometry, which is not well annotated by OSM tags. Therefore, unlike Task 1, which selects the OSM element with the largest size, Task 2 first identifies three OSM elements with the longest length and then select the one with the most tags as the most distinctive element to be described.
In the attribute interpretation step, we extract different attributes compared to Task 1 due to the geometrical differences between the ”area” and ”non-area” elements. The attributes for Task 2 are shown in Table 3. To better describe the position of the non-area element, we use endpoint locations instead of the coarse location derived from the element’s center. The nine-grid system, the same as used in Task 1, assigns labels to the endpoints. The sinuosity attribute describes the shape of the element, with categories including ”straight,” ”curved,” ”twisted,” ”closed,” and ”broken,” determined by thresholding the ratio between the element’s length and the distance between its endpoints. The normalized length attribute represents the relative length of the element compared to the square root of the image patch area. The length attribute is the actual length of the element in meters. The orientation attribute, determined by calculating the angle between the endpoints and the horizontal axis, describes the direction of the element with categories like ”W_E,” ”S_N,” ”SW_NE,” and ”NW_SE.” The simplified geometry attribute represents the simplified and normalized geometry of the element, following the same scheme as in Task 1. Finally, the is-cropped attribute indicates whether the element is cropped by the image patch.
Table 3: Attributes for distinctive non-area description in captioning Task 2
Attribute Example
Endpoint Locations ”(bottom-center, right-center)”, ”(left-bottom, right-bottom)”, etc.
Sinuosity ”straight”, ”curved”, ”twisted”, ”closed”, ”broken”
Normalized Length ”0.351”, ”1.260”, etc.
Length ”94”, ”339”, etc.
Orientation ”W_E”, ”S_N”, ”SW_NE”, ”NW_SE”
Simplified Geometry ”[(0.211, 0.000), (0.320, 0.794), …]”, etc.
Is-cropped True, False
The subsequent steps, tag interpretation and prompt assembly are largely similar between Task 1 and Task 2, with modifications in the prompt template to support the attributes specific to Task 2. To avoid redundancy, we do not elaborate on the details of these steps as they are consistent with those explained in Task 1. This structured approach ensures that each task is tailored to the unique characteristics of the elements being described, enhancing the accuracy and relevance of the LLM-generated captions for remote sensing data. The careful design of tasks and attributes enables a more effective analysis and utilization of the OSM data within the image patches.
2.2.2 Caption Augmentation Task
The caption augmentation task aims to enhance the diversity and quality of the initial captions generated for the image patches, creating multiple revisions from the raw caption with different tones. This task, denoted as Task 3 in this work and illustrated in Fig. 2, is relatively simple compared to previous tasks involving OSM element selection and interpretation, as it is purely a linguistic task and does not require specialized knowledge. Consequently, it comprises only two steps: revision example sampling and prompt assembly.
Unlike previous tasks that utilized a fixed example for each task, this task necessitates a larger number of examples to produce diverse and informative captions. To this end, we constructed a meta-example set containing multiple raw captions and their corresponding revisions. In this work, we separately selected five raw captions from Task 1 and Task 2 and generated five revisions for each raw caption, altering the tone, inflection, word choice, and length of the original caption. These revisions were written by human annotators and polished by GPT-4. In Task 3, examples are dynamically sampled from the meta-example set, selecting all five raw captions of the corresponding task and randomly choosing one of the five revisions for each raw caption. The five raw and revision pairs are then shuffled and used as the five-shot examples for the prompt assembly step. This approach ensures that the revisions generated by the LLM are diverse and informative.
In the prompt assembly step, a simpler prompt template is utilized, as shown in Fig. 4. The overall structure of the prompt template is similar to that of Tasks 1 and 2, divided into three parts: general instructions, few-shot examples, and task inputs. As the revision task is straightforward and does not require detailed caption objectives or specialized knowledge about the captioned element’s attributes, the general instructions are kept concise, informing the LLM to revise an existing caption, specifying the coordinate system used, and indicating the provided examples. The five-shot examples in the prompt template are defined by several placeholders, filled by sampled examples from the previous step. The task inputs include a placeholder for the caption to be revised, leaving the revised result empty for the LLM to fill in.
This method ensures that the final augmented captions are of high quality, diverse, and aligned with the intended use of the remote sensing data, thus facilitating the effective and insightful analysis.
Fig. 4: A brief showcase of the prompt template for Task 3. Details are replaced by ellipses for brevity. Placeholders are wrapped in braces and are to be filled dynamically by the task.
3 RSTeller Dataset
Utilizing the proposed dataset generation workflow, we have developed the RSTeller dataset. The data acquisition and processing were conducted using a local server equipped with 128 cores, 512 GB of memory, and four NVIDIA A800 GPUs. We deployed one Mixtral-7B instance on each GPU to generate the captions in parallel. The data acquisition process took approximately four months, during which we collected roughly 15 million image patches. Of these, approximately 7 million had corresponding OSM data fetched from the Kumi endpoint. The captioning process ran concurrently with data acquisition and processing, achieving a maximum throughput of 50,000 captions per day. This process spanned over two months and generated approximately 3 million captions. Ultimately, 2.5 million high-quality image-text pairs were selected and compiled into the RSTeller dataset. The following sections provide detailed insights into the dataset.
3.1 Dataset Overview
RSTeller dataset comprises approximately 1.2 million remote-sensing image patches, each accompanied by multiple captions, ranging from 2 to 5, resulting in roughly 2.5 million image-text pairs. The images are sourced from the National Agriculture Imagery Program (NAIP) on the GEE platform, with a ground sample distance (GSD) of 0.6 meters. Geographically, the dataset focuses on the United States, with images acquired via aircraft. Detailed attributes of the dataset are presented in Table 4.
Table 4: Attributes of the RSTeller dataset generated by the proposed workflow
Attribute Value
Number of Images 1,197,190
Number of Image-Text Pairs 2,539,256
Image Source National Agriculture Imagery Program (NAIP)
Ground Sample Distance (GSD) 0.6 meters
Pixel Dimensions $448\times 448$ pixels
Image Bands Red (R), Green (G), Blue (B)
Image Capture Date Range August 1, 2021 - November 26, 2022
Geographic Coverage United States
Image Acquisition Platform Aircraft
The distribution of captions within the dataset is illustrated in Fig. 5. As shown in Fig. 5(a), Task 2 captions slightly outnumber Task 1 captions, while Task 3 captions constitute more than half of the entire dataset. The prevalence of Task 3 captions is due to the multiple revisions of the raw captions. Fig. 5(b) indicates that while most image patches (88.44%) have two caption variations (one raw caption and one revision), some images have up to five variations, enhancing the dataset’s diversity in linguistic variations.
Fig. 5: Distribution of captions in the RSTeller dataset. (a) Proportions of captions categorized by tasks. (b) Relationship between the number of images and caption variations.
Fig. 6: Geographical distribution of caption density across the United States.
Rooted in NAIP, the dataset’s images cover the majority of the continental United States and parts of Hawaii. We surveyed the dataset and plotted the caption density by county, as shown in Fig. 6, to provide a better understanding of the dataset’s coverage and geographical distribution. Although the dataset covers most of the United States, some regions, such as Southern Mississippi and Northern Oregon, exhibit low or zero caption density. This is because the dataset was generated by sampling NAIP images based on their capture dates. For some regions, few NAIP images are available within the dataset’s date range (August 1, 2021, to November 26, 2022), resulting in low caption density.
To facilitate qualitative analysis, we have provided a sample inspection of the dataset in Fig. 7. This figure presents twenty image patches along with their corresponding OSM information and captions, sampled from various locations across the United States. The map shows red dots indicating the sample locations, with a digit labeling each patch ID. A topographic base map from OpenTopoMap [^8] https://opentopomap.org is included to provide a visual reference for the geographical locations of the samples. The figure then displays the image patches, selected OSM elements, their tags, and the captions under specific captioning tasks. Selected OSM elements are highlighted in red: area elements are highlighted with masks and enclosed by their respective boundaries, while non-area elements are highlighted with simple lines.
Overall, the RSTeller dataset is geographically and temporally diverse, covering a wide range of locations and time periods. It provides rich and varied semantics, making it suitable for training various VLMs for applications such as open-vocabulary classification, cross-modal retrieval, image captioning, and text-to-image generation in the remote sensing domain.
Fig. 7: Inspection of the RSTeller dataset.
Fig. 8: Inspection of the RSTeller dataset (continued).
3.2 Caption Analysis
This section evaluates the captions generated by the LLM to identify key attributes and potential biases within the RSTeller dataset. Our aim is to provide valuable insights that will inform the effective utilization of the dataset in future applications.
First, we examine the distribution of caption lengths, determined by counting the number of tokens using the Natural Language Toolkit (NLTK) word tokenizer. The caption lengths vary from 5 to 192 tokens, with a median length of 48 tokens and an average length of 54.2 tokens and its distribution is depicted in Fig. 9(a). This variation underscores the diversity within the captions, which could enhance the robustness of VLM training. However, given that the CLIP model has a maximum input length of 77 tokens, captions exceeding this limit may compromise performance, highlighting the need for meticulous preprocessing.
Moreover, we assess the metadata richness by analyzing the number of tags associated with each OSM element used for captioning. As Fig. 9(b) illustrates, while most images feature fewer than 10 tags, some are annotated with over 30, with a maximum of 63, indicating a skewed distribution with a long tail.
Fig. 9: Caption analysis of the RSTeller dataset. (a) Histogram of caption lengths. (b) Histogram of tag counts per image.
Subsequently, we scrutinize the most frequent non-stop words in the captions to assess content richness. The top 50 most frequent words are shown in Fig. 10. Common but meaningless stop words such as ”the”, ”is”, and ”and” are filtered out. The remaining words are sorted by their frequency in descending order. Words suggesting inferential reasoning like ”likely” and ”possibly” are prevalent, reflecting the LLM’s inference from the OSM tags and its knowledge. For example, in image patch 8 of Fig. 7, the OSM tag only contains the tag ”landuse: farmland,” but the caption generated by the LLM extends the semantics by adding ”It is likely used for growing crops,” connecting shallow tag information to deeper meanings. Additionally, words like ”meters,” ”within,” and ”orientation” appear frequently, demonstrating the richness of shape and geography information in the dataset due to the interpretation of attributes in the captioning tasks. For instance, the caption for image patch 3 in Fig. 7 precisely describes the road’s orientation (south to north) and length (270 meters).
Fig. 10: Top 50 most frequent words in the captions of the RSTeller dataset.
Additionally, we delve into the topical diversity of the captions by investigating the frequency of the top 20 keys and their corresponding values within the OSM tags used for captioning. As shown in Fig. 11(a), the top 20 most frequent keys include feature-related keys such as highway, natural, landuse, waterway, and building, while others are more general like name and TIGER (Topologically Integrated Geographic Encoding and Referencing), which are usually unique identifiers. To further explore the attribute distribution of some features, we survey the top 20 most frequent values from keys highly related to image features. Fig. 11(b-f) display these for highway, natural, landuse, waterway, and surface. All values exhibit long-tailed distributions, with significant differences in frequency. Notably, even the most frequent key ”highway” has a frequency of no more than 500k, which is less than half of the total number of OSM elements used for captioning the entire dataset, indicating a diverse and rich set of attributes used for captioning.
Fig. 11: Topic distribution within the captions of the RSTeller dataset. (a) Top 20 most common OSM keys used in captions. (b-f) Top 20 most frequent values for selected OSM keys used in captioning.
In conclusion, while the captions in the RSTeller dataset are rich and diversified, focusing on varied remote sensing features, they also contain ambiguities and inferred content that may introduce noise. These attributes are pivotal for training sophisticated VLMs in the remote sensing domain, but they also necessitate careful handling to mitigate potential biases and ensure the utility of the trained models.
4 Experiments
To validate the proposed automated dataset generation workflow and assess the effectiveness of the RSTeller dataset, we performed a comprehensive series of experiments. These experiments involved the continual pre-training of the CLIP model with the RSTeller dataset, followed by evaluations of its performance on zero-shot image classification and retrieval tasks. The purpose of these experiments was to rigorously test several hypotheses regarding the utility and impact of the dataset on the training of robust VLMs tailored for remote sensing applications.
The experimental series is structured as follows:
  1. 1.
    Data Scalability and Model Performance Analysis: The first experiment explores the relationship between the performance of VLMs in remote sensing and the scale of domain-specific data, specifically examining the influence of the number of RS-related image-text pairs. This experiment aims to understand how data volume affects model accuracy and generalization capabilities in open-set conditions.
  2. 2.
    Dataset Effectiveness Evaluation: We then evaluate the effectiveness of the RSTeller dataset through continual pre-training using various CLIP model checkpoints. This step assesses whether incremental learning with our dataset can enhance the model’s performance on standard benchmarks. Additionally, this experiment explores the impact of different pre-trained checkpoints on the model’s performance.
4.1 Data
The training data utilized in our experiments consists of three distinct groups: the RSTeller dataset, the SkyScript (Wang et al., 2024) dataset, and the LAION-10M dataset, which is a subset of the LAION-400M (Schuhmann et al., 2021). The detailed information is provided in Table 5 and explained below.
Table 5: Datasets utilized for training
Domain Dataset Image-Text Pairs
RS RSTeller 2,539,256
SkyScript 2,554,817
Common LAION-10M 10,003,420
The RSTeller dataset, which we propose, consists of approximately 2.5 million image-text pairs. There are more than two captions for one image patch in this dataset. To ensure clarity and prevent redundancy during the training of the CLIP model, each image patch appears only once in a training batch. We implement a sampling strategy where image patches for training batches are selected first, followed by a random selection of captions for these images. Both raw and augmented captions have an equal probability of 50% of being chosen for training.
The SkyScript dataset, similarly sourced from OSM and GEE, enriches our training with an additional 2.5 million image-text pairs. Each image in this dataset is uniquely associated with one caption. This dataset is integral to augmenting the RS domain knowledge and is merged with the RSTeller dataset to foster a robust study of the relation between model performance and domain data scale in our first experiment.
Lastly, the LAION-10M dataset represents a selective download from the LAION-400M dataset’s metadata. Despite some inaccessible URLs, approximately 10 million image-text pairs were compiled. This dataset functions as a supplementary component in our experiments, providing a diverse common knowledge base that aids in preventing catastrophic forgetting and reduces the risk of the model overfitting to specific domain data. This dataset is used across all the experiments as a regulatory fraction of model training.
For evaluative purposes, we employ a series of zero-shot classification and image retrieval tasks. For zero-shot classification, we use seven widely recognized and comparatively extensive benchmark datasets, including AID (Xia et al., 2017), EuroSAT (Helber et al., 2019), fMoW (Christie et al., 2018), Million-AID (Long et al., 2021), PatternNet (Zhou et al., 2018), NWPU-RESISC45 (RESISC) (Cheng et al., 2017), and RSI-CB256 (RSI-CB) (Li et al., 2020). For image retrieval evaluation, three common benchmark datasets are used, including UCM Captions (Qu et al., 2016), RSICD (Lu et al., 2017), and RSITMD (Yuan et al., 2022). Details of these benchmark datasets are presented in Tables 6 and 7, following the data usage protocols established in the prior study as SkyScript.
Table 6: Benchmark datasets for zero-shot classification evaluation
Dataset Split Classes Images
AID test 30 2,000
EuroSAT test 10 2,700
fMoW val 62 106,081
Million-AID train 51 10,000
PatternNet train 38 30,400
RESISC train 45 31,500
RSI-CB train 35 24,747
Table 7: Benchmark datasets for zero-shot image retrieval evaluation
Dataset Split Captions Images
UCM Captions test 1,050 210
RSICD test 5,465 1,093
RSITMD test 2,260 452
4.2 CLIP Continual Pre-training
VLMs are typically pre-trained on a large, predefined dataset and have their parameters fixed upon the completion of pre-training. However, given the dynamic nature of the world, the static pre-training dataset may not adequately encompass the knowledge required for specific downstream tasks. To address this gap, continual pre-training is employed to integrate new knowledge into an existing model by incorporating domain-specific data and adhering to the established pre-training paradigm.
Our study adopts continual pre-training to investigate the relationship between the scale of domain-specific data and performance enhancements in the CLIP model. We examine the influence of our innovative RSTeller dataset on model performance. Known challenges associated with training exclusively on new domain data include significant performance degradation due to catastrophic forgetting and potential overfitting. To mitigate these challenges, our experiments incorporate data from the original domains, specifically LAION-10M, throughout the continual pre-training process. We consistently sample one-third of the data from LAION-10M for each training batch while the rest are sampled from the RS domain. This technique is applied across all our experiments to ensure balanced training and maintain model stability.
We analyzed two versions of CLIP models: ViT-B/32 and ViT-L/14. The ViT-B configuration utilizes 12 transformer layers and 12 attention heads, whereas the ViT-L configuration comprises 24 transformer layers and 16 attention heads. The denominations /32 and /14 represent the respective patch sizes, which are $32\times 32$ and $14\times 14$ pixels, processed by the model tokenizer. Initial model setups are derived from various checkpoints, including those from the OpenCLIP project (Ilharco et al., 2021) trained on the LAION-2B dataset (Schuhmann et al., 2022), which provides a baseline of regularly pre-trained models on generic web-based image-text data. Additionally, we utilized checkpoints commonly employed by the industrial community, trained on LAION-400M (Schuhmann et al., 2021), WIT (Radford et al., 2021), and DataComp (Gadre et al., 2023), to ascertain the benefits conferred by our RSTeller dataset in the RS domain. A brief comparison of the data amount among these datasets and all the checkpoints used in our experiments are presented in Table 8. The model checkpoints are represented by the identifier used in the OpenCLIP repository (Ilharco et al., 2021).
Table 8: Pre-training datasets and pre-trained checkpoints
Dataset Size ViT-B/32 ViT-L/14
WIT 0.4B openai openai
LAION-400M 0.4B laion400m_e32 laion400m_e32
DataComp 1.4B datacomp_xl_s13b_b90k datacomp_xl_s13b_b90k
LAION-2B 2.3B laion2b_s34b_b79k laion2b_s32b_b82k
Fig. 12: Relationship between domain data scale and zero-shot classification performance. Left: A ViT-B/32 model. Right: A ViT-L/14 CLIP model. Both models are initialized with weight trained on LAION-2B. Scaling domain data consistently improves the performance of the model on the validation dataset.
Fig. 13: Relationship between domain data scale and zero-shot image retrieval performance. Left: A ViT-B/32 model. Right: A ViT-L/14 CLIP model. Both models are initialized with weight trained on LAION-2B. Scaling domain data consistently improves the performance of the model on the validation dataset.
Experimentation was conducted using the OpenCLIP framework, incorporating automatic mixed-precision (AMP) to expedite training and optimize performance. For the experiments that scale domain data (Section 4.3), gradient checkpointing was employed to minimize memory usage, and the standard Adam optimizer was utilized with a learning rate of $1\times 10^{-5}$. Both linear warm-up and cosine annealing learning rate schedules were implemented. Each experiment involved random iterations over the data for 5,000 steps with a total batch size of 8,192 across 4 NVIDIA A800 GPUs. For the dataset effectiveness evaluation experiments (Section 4.4), gradient checkpointing was disabled, and a total batch size of 1,024 was used on 4 NVIDIA A800 GPUs for more efficient training. The training steps and learning rate were adjusted to 20,000 steps and a learning rate of $1\times 10^{-8}$, respectively, to optimize performance.
4.3 Model Performance wrt. Domain Data Scale
This experiment was designed to investigate whether scaling the domain-specific data enhances model performance. To this end, we amalgamated our RSTeller dataset with the SkyScript dataset to create a robust RS domain dataset comprising approximately 5 million image-text pairs. This extensive dataset allows for a detailed analysis of the model’s performance as it relates to variations in domain data volume. We partitioned the data into four equal subsets, each representing about one-fourth of the total data, and ensured no overlap between the subsets. In our experiments, we incrementally incorporated these subsets into the training process, thereby systematically increasing the domain data scale. Models pre-trained on LAION-2B were used as the starting point for the continual pre-training. The models were then trained on approximately 25%, 50%, 75%, and 100% of the domain data, respectively.
The results for zero-shot classification are shown in Fig. 12. The x-axis represents the volume of RS-related domain data used for continual pre-training, while the y-axis indicates the top-1 classification error on the validation dataset. Initially, we evaluated the loaded checkpoints directly on benchmark datasets to establish a baseline for models pre-trained on generic web-based image-text data, represented as stars in the figure. Subsequently, we continually pre-trained the models on RS domain-specific data and evaluated them on the same benchmarks, with the results denoted as dots in the figures. To better visualize performance trends, linear fits were applied to the data points, shown as dashed lines, with the slope values $k$ annotated to reflect the changes in error rates.
Most benchmarks show improvement when domain data is scaled up. Notably, the slopes of both models’ error rates are negative ($k_{Average}=-0.75$ for ViT-B/32 and $k_{Average}=-0.34$ for ViT-L/14). Benchmarks such as PatternNet, RESISC, and AID exhibit substantial improvements, as indicated by the slopes of their linear fits (e.g., $k_{PatternNet}=-1.63$, $k_{RESISC}=-1.03$ for the ViT-B/32 model, and $k_{AID}=-1.36$ for the ViT-L/14 model). These results underscore the positive impact of enlarging the domain-specific dataset on the model’s classification accuracy. The EuroSAT on the ViT-L/14 model, on the other hand, demonstrates an unexpected increase in top-1 classification error as the domain data scale rises ($k_{EuroSAT}=0.76$). This anomaly suggests potential noise in the training process or an inherent difficulty in this benchmark that requires more careful consideration in scaling the domain data.
For some benchmarks, such as AID and RSI-CB, there is an initial performance degradation when domain-specific data is first introduced. In extreme cases, even with all the domain data used for training, the fine-tuned model performance still does not surpass the baseline (e.g., AID for ViT-B/32). Nevertheless, the continual inclusion of domain data leads to consistent performance improvement. For instance, the AID benchmark shows a steady decline in error rates despite the initial degradation, as indicated by the slopes ($k_{AID}=-0.72$ for ViT-B/32 and $k_{AID}=-1.36$ for ViT-L/14). This trend suggests that further scaling of domain data could potentially surpass the baseline performance of models pre-trained on generic web-based data.
The varied slopes of the linear fits also highlight the differential impact of domain data across different benchmarks. For example, in the ViT-B/32 experiment, the significant negative slope observed for PatternNet ($k_{PatternNet}=-1.63$) compared to the minimal impact on Million-AID ($k_{Million-AID}=-0.07$) suggests that the effectiveness of domain-specific data augmentation varies across different datasets. This variation indicates inherent differences in the benchmarks themselves and possibly the domain specificity of the data they contain.
The experimental results for zero-shot image retrieval tasks, illustrated in Fig. 13, reveal trends consistent with those observed in the classification experiments. The x-axis represents the volume of RS-related domain data used for continual pre-training, while the y-axis indicates the evaluation metric, which is the average recall rates of top-1 (R@1), top-5 (R@5), and top-10 (R@10), subtracted from 100.
For both the ViT-B/32 and ViT-L/14 models, there is a general improvement in performance with increasing domain data, as indicated by the negative slopes of the linear fits ($k_{Average}=-0.33$ for ViT-B/32 and $k_{Average}=-0.57$ for ViT-L/14). This trend is evident across multiple benchmarks, highlighting the benefits of enlarging domain-specific datasets for image retrieval tasks.
However, a notable difference from the classification results is observed in the ViT-L/14 model’s performance for the retrieval task. None of the fine-tuned ViT-L/14 models were able to surpass the original checkpoint performance across all benchmarks, even with the full domain data included. Despite this, there is a consistent trend of improving recall rates with the increase in domain-specific data, suggesting that further scaling could potentially lead to superior performance in the future.
The performance of the ViT-L/14 model, both on classification and retrieval tasks, whether fine-tuned or not, consistently outperforms the ViT-B/32 model in most benchmarks. This is apparent from the lower error rates and higher recall rates across both the original model checkpoints and the models fine-tuned on all the domain data, indicating the superior capability of a larger model in handling various tasks.
The experimental results confirm that increasing domain-specific data generally enhances model performance in the field of remote sensing. Despite initial degradation observed in specific benchmarks, continuous improvement with increased data volume underscores the potential for further advancements. These findings validate the significance of this work, highlighting the proposed automated workflow’s efficiency in generating larger datasets at a low cost. This capability offers substantial benefits to the research community, particularly for training VLMs related to remote sensing.
4.4 Dataset Effectiveness
Table 9: Zero-shot top-1 classification accuracy (%) on benchmark datasets before and after continual pre-training with RSTeller dataset
Model Pretrained Dataset AID EuroSAT fMoW Million-AID RESISC PatternNet RSI-CB Average
ViT-B/32 LAION-2B 68.10 24.40 16.00 53.75 60.46 64.36 38.11 46.45
+RSTeller 61.15 31.11 14.70 51.36 54.36 66.37 40.85 45.70
$\Delta$6.95 6.71 -1.30 -2.39 -6.10 2.01 2.74 -0.75
LAION-400M 60.20 29.19 13.12 43.91 52.50 51.02 31.85 40.26
+RSTeller 60.25 37.93 14.25 41.14 47.36 50.37 36.90 41.17
$\Delta$ 0.05 8.74 1.12 -2.77 -5.14 -0.66 5.05 0.91
WIT 57.40 25.04 15.42 51.52 54.44 52.80 29.92 40.93
+RSTeller 59.65 36.44 15.60 52.79 55.13 62.29 35.83 45.39
$\Delta$ 2.25 11.41 0.18 1.27 0.69 9.48 5.92 4.46
DataComp 66.10 24.15 15.47 49.47 62.56 60.14 35.24 44.73
+RSTeller 63.65 36.89 15.61 49.73 58.70 64.28 41.39 47.18
$\Delta$2.45 12.74 0.14 0.26 -3.86 4.13 6.15 2.44
ViT-L/14 LAION-2B 67.80 38.04 24.39 63.02 70.42 72.79 42.92 54.20
+RSTeller 66.05 45.33 21.29 60.96 68.54 76.22 46.23 54.95
$\Delta$1.75 7.30 -3.10 -2.06 -1.88 3.42 3.31 0.75
LAION-400M 66.55 33.67 19.85 59.72 66.11 67.11 43.62 50.95
+RSTeller 65.05 47.07 17.36 56.89 62.14 69.83 41.89 51.46
$\Delta$1.50 13.41 -2.49 -2.83 -3.97 2.72 -1.73 0.52
WIT 69.05 41.70 25.48 57.72 66.62 71.32 43.05 53.56
+RSTeller 67.60 56.30 22.30 58.16 64.68 77.14 47.32 56.21
$\Delta$1.45 14.59 -3.18 0.44 -1.95 5.82 4.27 2.65
DataComp 69.65 52.56 27.57 66.89 70.98 78.22 43.59 58.49
+RSTeller 68.85 55.22 23.15 65.24 69.96 79.82 47.71 58.57
$\Delta$0.80 2.67 -4.42 -1.65 -1.02 1.60 4.13 0.07
To assess the benefits of the RSTeller dataset for pre-trained models, a series of continual training experiments were conducted on various CLIP checkpoints, evaluating their zero-shot capabilities before and after the continual pre-training. Notably, in this experiment, the SkyScript data was excluded from the training data, differing from the previous experiment. The LAION-10M was also utilized to prevent catastrophic forgetting and overfitting.
The evaluation results are presented in Tables 9 and 10. These tables report the top-1 accuracy for zero-shot classification tasks and the average recall rate for zero-shot text-to-image retrieval tasks. For each pre-trained checkpoint, the performance of the baseline checkpoints on the benchmark datasets, the performance after continual pre-training on the RSTeller dataset, and the corresponding improvements (marked as $\Delta$) are shown. The positive improvements are highlighted in red, while the negative changes are highlighted in blue. For each benchmark, the largest and least improvements are highlighted in boldface, and the best results within the continual pre-trained models are underlined.
Table 9 summarizes the top-1 zero-shot accuracy on various benchmark datasets before and after continual pre-training on the RSTeller dataset. Except for the ViT-B/32 model pre-trained on LAION-2B, all other models show improvements in the average results (marked as $\Delta$ values in red). This suggests that the RSTeller dataset enhances the performance of the VLMs, making it a valuable resource for pre-training VLMs in the remote sensing domain.
Despite the overall improvements in the average zero-shot classification results, the degree of enhancement varies across different benchmarks. EuroSAT consistently shows the most significant improvements (marked as bold $\Delta$ values in red) for all models and pre-trained checkpoints, with a maximum increase of $14.59\%$ for the ViT-L/14 model pre-trained on WIT. In contrast, some benchmark datasets such as AID, fMoW, and RESISC experienced notable decreases. However, these decreases were not consistent across all pre-trained checkpoints. For example, the ViT-B/32 model pre-trained on LAION-2B showed a significant decrease in AID by $6.95\%$, whereas the same model pre-trained on WIT saw an increase of $2.25\%$ on the same benchmark. This variability suggests that the RSTeller dataset may favor certain pre-trained checkpoints and tasks.
Continual pre-training from different checkpoints for the same model also produces varying results. For both ViT-B/32 and ViT-L/14 models, checkpoints from DataComp consistently demonstrated advantages over other pre-trained models, ranking highest with respect to the average accuracy across seven benchmarks. For most individual benchmarks, checkpoints from DataComp also achieved many of the top results (values with underlines). These findings indicate that a good starting checkpoint for continual pre-training can be advantageous, and checkpoints trained on DataComp are strong candidates for continual pre-training tasks in the RS domain.
Experiment results on zero-shot image retrieval show similar outcomes. As demonstrated in Table 10, most checkpoints benefit from continual pre-training on the RSTeller dataset across the benchmark datasets. Notably, there is a slight difference between classification and retrieval tasks: while the checkpoint from DataComp outperforms others for the ViT-L/14 model, the checkpoint from LAION-2B achieves the best results for the ViT-B/32 model after continual pre-training. This variation highlights the nuanced differences between classification and retrieval tasks.
Table 10: Zero-shot image-to-text retrieval recall rates (%) on benchmark datasets before and after continual pre-training with RSTeller dataset
Model Pretrained Dataset RSICD RSITMD UCM-Captions Average
ViT-B/32 LAION-2B 16.73 27.50 57.21 33.81
+RSTeller 17.28 27.59 54.55 33.14
$\Delta$ 0.55 0.09 -2.65 -0.67
LAION-400M 14.49 22.54 54.38 30.47
+RSTeller 15.40 24.43 53.93 31.25
$\Delta$ 0.91 1.89 -0.44 0.78
WIT 12.24 22.34 42.44 25.67
+RSTeller 16.77 27.51 51.37 31.89
$\Delta$ 4.53 5.18 8.93 6.21
DataComp 15.88 25.03 52.87 31.26
+RSTeller 15.10 26.49 53.40 31.67
$\Delta$0.78 1.46 0.53 0.41
ViT-L/14 LAION-2B 20.34 31.81 62.95 38.37
+RSTeller 20.66 32.91 60.92 38.16
$\Delta$ 0.32 1.10 -2.03 -0.21
LAION-400M 16.18 26.60 54.47 32.41
+RSTeller 16.49 28.52 53.76 32.92
$\Delta$ 0.32 1.92 -0.71 0.51
WIT 13.81 24.08 56.76 31.55
+RSTeller 17.42 28.99 57.82 34.75
$\Delta$ 3.61 4.91 1.06 3.19
DataComp 21.04 31.01 59.59 37.21
+RSTeller 21.27 31.32 61.98 38.19
$\Delta$ 0.22 0.31 2.39 0.97
In conclusion, the RSTeller dataset offers researchers a robust starting point to expand their datasets, facilitating improved training of their own VLMs. The positive impact observed across multiple benchmarks underscores the dataset’s potential to enhance model performance in the remote sensing domain, thus making it a valuable asset for future research and practical applications.
5 Discussions
The above experimental results provide several key insights into the impact of scaling domain-specific data on the performance of VLMs as well as the continual pre-training in the remote sensing domain. The findings offer valuable guidance for future research and practical applications in dataset generation and model training.
First, the results in domain data scaling (Section 4.3) confirm that increasing the volume of domain-specific data generally enhances the model performance across various benchmarks. This is evidenced by the consistent improvement in both zero-shot classification and image retrieval tasks as more domain data is included. However, the degree of improvement varies among different benchmarks, indicating that the effectiveness of domain-specific data augmentation is influenced by the nature of the benchmark dataset. For instance, benchmarks like PatternNet and AID showed significant performance gains, whereas others like EuroSAT exhibited less pronounced improvements or even anomalies in certain cases.
One notable observation is that the initial performance degradation observed in some benchmarks, such as the AID and RSI-CB, when domain-specific data is first introduced. Despite this initial setback, continuous scaling of the domain data eventually leads to performance improvements, suggesting that the models require a certain threshold of domain-specific data to effectively learn and generalize. This indicates the importance of not being discouraged by the initial performance drops and continuing to expand the dataset for long-term gains.
The ViT-L/14 model consistently outperformed the ViT-B/32 model across most benchmarks, regardless of whether the models were fine-tuned or not. This underscores the advantage of larger models in handling diverse and complex tasks in the RS domain. However, the fine-tuned ViT-L/14 models failed to surpass the original checkpoint performance in some of the benchmarks, even with the full domain data included. This suggests that the overall distribution of the domain dataset created in this experiment could still be different from the benchmark datasets, causing a substantial domain gap. Increasing the volume of data from more diverse sources may bridge this gap.
Finally, the experiments in testing the RSTeller dataset (Section 4.3) revealed that the starting checkpoint plays a critical role in the effectiveness of continual pre-training. Models pre-trained on DataComp outperformed those pre-trained on other datasets in most cases, suggesting that selecting a robust initial checkpoint can lead to better results in continual pre-training. This finding emphasizes the importance of choosing the right pre-trained model as a starting point for further training with the RSTeller dataset.
Based on these findings, several recommendations can be made for future works when trying to bring the VLMs to the RS domain:
Diverse Data Sources: When generating RS-domain datasets, it is essential to include data from a variety of sources. This diversity can help the model learn more robust features and improve generalization across different benchmarks. For example, incorporating images with varying ground sample distances, spectral bands, and geographic locations can enhance the dataset’s comprehensiveness and utility.
Threshold Data Volume: Researchers should aim to exceed a certain threshold of RS-domain data volume to overcome the initial performance degradation and achieve substantial improvements. This highlights the need for large-scale data collection efforts and the importance of not halting the process prematurely.
Complex Models: Utilizing more complex models, such as ViT-L/14, can lead to better performance in handling varied tasks within the RS domain. Future research should explore even larger and more sophisticated models that can leverage the enriched domain-specific data effectively.
Long-Term Data Scaling: The continuous improvement observed with increasing data volume suggests that long-term efforts in scaling domain-specific datasets will yield better results. Researchers should consider ongoing data collection and integration as a key strategy for advancing the field.
Selection of Initial Checkpoints: If choosing to build a VLM from existing pre-trained models, it is essential to identify and utilize robust initial checkpoints that can maximize the benefits of continual pre-training with the RS domain data.
6 Limitations of the Proposed Work
Despite its strengths, our approach that scales RS domain multimodal data for training RS VLMs has the following limitations that warrant future investigation.
First, the captioning tasks within our dataset generation workflow are simplistic, typically focusing on a single element, which might not capture all pertinent details in the images. For instance, image patch 4 in Fig. 7 is captioned solely by its depiction of an industrial complex, omitting other significant elements like cars and the parking lot within. Similarly, image patch 10 focuses only on a single solar panel within a larger solar power plant, potentially misleading in terms of scale, location, and quantity. Future enhancements could include more intricate captioning tasks to encompass a broader range of information within the images.
Moreover, while the RSTeller dataset is a valuable resource, its scale and diversity are not yet comparable to larger datasets such as LAION-400M. It is sourced from a single platform with a fixed ground sample distance and specific image bands, and its geographic coverage is limited to the United States. Subsequent efforts should aim at improving the dataset’s quality and expanding its scale to encompass a more diverse array of geographies, image bands and so on.
Finally, the VLM analyzed in this study, which is a relatively simple CLIP model, does not incorporate recent advancements that would enable more sophisticated interactions with humans in natural language. Future research should explore more complex models that can better understand and utilize the semantics of captions and interact more naturally with human users.
7 Conclusion
In this paper, we have meticulously developed and presented an automated dataset generation workflow that enables researchers to efficiently gather and utilize multimodal data for RS applications on a large scale. This workflow optimizes the use of existing, openly available remote-sensing data and employs an LLM to transform raw data into a high-quality, usable dataset for training VLMs in the remote sensing domain. Central to our workflow is the introduction of RSTeller, a multimodal dataset comprising over 2.5 million high-quality image-text pairs with diverse and informative captions, suitable for a wide range of research tasks. Utilizing this dataset for training RS VLMs not only facilitates access to high-quality data for researchers but also reduces the environmental impact associated with generating and maintaining such a dataset. Moreover, we highlight the potential for significant advancements in the field of RS VLMs through the continual enrichment of RS-related multimodal data. We envision future research expanding these datasets to a larger scale, thus enabling the training of more robust VLMs and enhancing the potential prosperity achievable in this field.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
During the preparation of this work the authors used ChatGPT in order to improve readability and language. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication
Acknowledgments
This study was supported by the National Natural Science Foundation of China (Nos. 62101416, 62301405), the Natural Science Basic Research Program of Shaanxi, China (2022JQ-708), and Fundamental Research Funds for the Central Universities, China.